jhonyy9 (jhonyy9):

- from this equation how many will be ,,b" ? 1= -2b(lg((x+yb)/(ab))

OpenStudy (dumbcow):

could you clarify the question

jhonyy9 (jhonyy9):

b= ?

jhonyy9 (jhonyy9):

lg is logarithm on base 10

OpenStudy (dumbcow):

\[\frac{\log(x+by)}{ab} = -\frac{1}{2b}\] \[\log(x+by) = -\frac{a}{2}\] \[x+by = 10^{-a/2}\] \[b = \frac{10^{-a/2}-x}{y}\]

jhonyy9 (jhonyy9):

thank you but there is lg( (x+by)/(ab) )

jhonyy9 (jhonyy9):

can you solve it ,help me if there is so ?